Sweet Inc. uses a calendar year for financial reporting. The company is authorized to issue 9,720,000 shares of $10 par common stock. At no time has Sweet issued any potentially dilutive securities. Listed below is a summary of Sweet’s common stock activities.

1. Number of common shares issued and outstanding at December 31, 2018

1,910,000

2. Shares issued as a result of a 10% stock dividend on September 30, 2019

191,000

3. Shares issued for cash on March 31, 2020

1,830,000

Number of common shares issued and outstanding at December 31, 2020

3,931,000

4. A 2-for-1 stock split of Sweet’s common stock took place on March 31, 2021

Compute the weighted-average number of common shares used in computing earnings per common share for 2019 on the 2020 comparative income statement.

shares

Compute the weighted-average number of common shares used in computing earnings per common share for 2020 on the 2020 comparative income statement.

shares


Compute the weighted-average number of common shares to be used in computing earnings per common share for 2020 on the 2021 comparative income statement.

shares

Compute the weighted-average number of common shares to be used in computing earnings per common share for 2021 on the 2021 comparative income statement.

shares
Answer #1

Please find below table useful to compute desired results: -

9720000 1910000 =10%*B4 1830000 =SUM(B4:36) =B4*9/12 =1.1 =B9*B10 =(B4+B5)*3/12 =B11+B12 А 1 Particulars given: 2 Authorized

End results would be as follows: -

A 9,720,000 10 1,910,000 191,000 | 1.830.000 3,931,000 1,432,500 1.1 1,575,750 525,250 2,101,000 1 Particulars given 2 Author